Use 0, 1, 2, 3, 4 to substitute for x in the formula x2-x+41; see if prime number occurs then find a number for x that when substituted in the formula yields a composite number.




Answer by Edwin McCravy(20060)

Use 0, 1, 2, 3, 4 to substitute for x in the formula x²-x+41; see if prime number occurs then find a number for x that when substituted in the formula yields a composite number.

x    x²- x + 41         prime?
0    0²- 0 + 41 = 41     yes
1    1²- 1 + 41 = 41     yes
2    2²- 2 + 41 = 43     yes
3    3²- 3 + 41 = 47     yes
4    4²- 4 + 41 = 53     yes


It's obvious that when 41 is substituted for
x, the middle term will cancel the last
term leaving 41² which is composite. 

So when x = 41, x²- x + 41 = 41²- 41 + 41 = 41² = 1681, not prime.

Edwin